# Creativity and Machine Learning: A Survey

### v4

@@ 3.2.3. Applications
GANs have been applied to a variety of practical problems in several application scenarios. They have been widely used for semi-supervised learning (Odena, 2016); for generating adversarial examples (Xiao et al., 2018) to better train image classifiers (Madry et al., 2018); and, in general, in computer vision (see (Wang et al., 2021) for a detailed discussion). The generative power of GANs has also found its place in recommender systems (see (Deldjoo et al., 2021)) to generate fashion items; in science and chemistry (Mendez-Lucio et al., 2020; Motamed et al., 2021). Of course, its ability to generate high-quality samples has been exploited in many other areas, from anime design (Jin et al., 2017) and 3D object modeling (Wu et al., 2016) to photo-realistic consequences of climate change (Schmidt et al., 2022). Conditional inputs also allow to produce artistic works by controlling stylistic properties such as genre (Tan et al., 2017) or influencer (Chu, 2018).
Finally, the most famous example of the artistic power of GAN is the collection of paintings by Obvious, a French art collective (Vernier et al., 2020); one of their works has been sold to more than 400,000 dollars [^3] Fun fact: the sold painting is called Portrait of Edmond De Belamy because Belamy sounds like bel ami, a sort of French translation of… Goodfellow..
@@ 3.3.1. Core Concepts
A sequence prediction model is a generative model that considers generation as a sequential process. It works in an autoregressive fashion: it predicts the future outcome of the sequence (i.e., the next token) from the previously observed outcomes of that sequence, usually by means of an internal state that encodes information from the past. It is trained to minimize the prediction error for each token in the dataset. At inference time, this simple yet effective approach only requires to produce one token after the other, feeding back to the model what has been produced so far (Karpathy, 2015). It makes possible to learn dependencies between tokens in real data, so that the same dependencies can be exploited when generating synthetic data. However, this causes the generation to be highly dependent from real data, e.g., there is the risk of potentially reproducing portions of the training set.
@@ 3.3.2. Examples of Models
Sequence prediction models are also used for domains not commonly modeled as sequences, like images. Image modeling can be defined in a discrete way by means of a joint distribution of pixels: the model learns to predict the next pixel given all the previously generated ones. It starts at the top left pixel and then proceeds towards the bottom right. The two seminal architectures for sequence prediction of images are PixelRNN and PixelCNN (Van Den Oord et al., 2016b). The former is a two-dimensional RNN (based on rows or on diagonals). The latter is a convolutional neural network (CNN) with an additional fixed dependency range (i.e., the convolution filters are masked in order to only use information about pixels above and to the left of the current one). To obtain better results, gated activation units can be used in place of rectified linear units between the masked convolutions; conditional inputs encoding high-level image descriptions can be used as well (Van Den Oord et al., 2016c). Notably, the Gated PixelCNN architecture can also be used for other types of data: WaveNet (Van Den Oord et al., 2016a) implements it to generate audio based on the waveform, possibly guiding the generation with conditional inputs.
@@ 3.3.3. Applications
As discussed, sequence prediction models have been used to learn to write poems or stories (by predicting a character, syllable, or word after the other); to compose music (by predicting a note or a waveform after the other); to draw images (by predicting a pixel after the other). In general, they can be used for any kind of time series forecasting (Lim and Zohren, 2021). They can also be used for co-creativity, as in Creative Help (Roemmele and Gordon, 2018).
Despite their simplicity, sequence prediction models are one of the most successful generative techniques. An interesting example is Sunspring. It might be considered as the first AI-scripted movie: it was generated by a Char-RNN trained on thousands of sci-fi scripts (Miller, 2019). The quality of the result is demonstrated by the fact that it was able to reach the top ten at the annual Sci-Fi London Film Festival in its 48-Hour Film Challenge [^4] Quite interestingly, the AI that wrote Sunspring declared that its name was Benjamin, probably in honor of Walter Benjamin, the German philosopher who, already in 1935 (Benjamin, 2008), understood that new mechanical techniques related to art can radically change the public attitude to art and artists..
@@ 3.3.4. Critical Discussion
Sequence prediction models generate outputs that have characteristics of both exploratory and combinatorial creativity. They are based on probabilistic predictions and they are able to generate new outputs in the induced space, but they can also reuse sequences of tokens from different works, combining them together. There is no guarantee that the results will be valuable or novel; and classic methods such as RNNs lack surprise (Bunescu and Uduehi, 2019). It is worth noting that the possibility of using conditional inputs and being able to work at different levels of abstraction might indirectly lead to creative outputs, but creativity should then be attributed to the higher-level component (or human, if the input is provided by the user) that is guiding the generation for specific elements and characteristics of the result.
@@ 3.4.1. Core Concepts
In principle, a Transformer is nothing more than an autoregressive model: it works by predicting the current token given the previous ones (see Section 3.3). However, few fundamental differences exist. A Transformer can also be trained by means of masked modeling: some of the input tokens are randomly masked, and the model has to learn how to reconstruct them from the entire context, and not only from the previous portions (Devlin et al., 2019). The possibility of dealing with very long sequences allows for prompting. By providing a natural language prompt in input, the model is able to generate the desired output, e.g., the answer to a question, a class between a set of classes for a given text, or a poem in a particular style (Brown et al., 2020). This is done by simply passing the prompt in input as a text, and then leveraging the model to predict what comes next (e.g., the answer to a question).
These advantages, together with the massive amount of data available, the increasing computational power, and the parallelism induced by their architecture, have led Transformer-based models to become the state of the art for several tasks. Nevertheless, the computational costs of the architecture from (Vaswani et al., 2017) grow quadratically with the input size.
@@ 3.4.2. Examples of Models
The domain mostly influenced by Transformers is natural language processing (NLP). Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2019) is a Transformer-based encoder trained for both predicting the next sentence (in an autoregressive fashion) and reconstructing masked tokens from the context. Several enhanced variations of the original model have been proposed, such as, for instance, solutions that remove the next-sentence pre-training objective (Liu et al., 2019), use inter-sentence coherence as an additional loss (Lan et al., 2020), or employ distillation (Hinton et al., 2015) to train a smaller model (Sanh et al., 2019). The other main approach is that used by the Generative Pre-trained Transformer (GPT) family (Radford, 2018; Radford et al., 2019; Brown et al., 2020). Here, a Transformer-based decoder is trained in an autoregressive way by additionally conditioning on the task of interest. After training, it can be used to perform a wide range of tasks by providing a description or few demonstrations of the task. The effectiveness of this text-to-text generative approach has then been explored by T5 (Raffel et al., 2020). Many other large language models (Shoeybi et al., 2019; Touvron et al., 2023a; Zhang et al., 2022) have been proposed to achieve better results by means of more parameters and computation (Smith et al., 2022), or more qualitative data (Gunasekar et al., 2023). Mixture of Experts (Shazeer et al., 2017) can be used as well in place of feed-forward network to train a larger but lighter model (since only portions of it are used per task), as done by Generalist Language Model (GLaM) (Du et al., 2022). Finally, Bidirectional and Auto-Regressive Transformer (BART) (Lewis et al., 2020) ideally merges together a BERT-encoder (trained by corrupting text with an arbitrary noising function) and a GPT-decoder (trained to reconstruct the original text autoregressively). Such an encoder-decoder architecture is able to achieve state-of-the-art results in machine translations, as well as other text-to-text tasks.
@@
Transformers have been highly influential in computer vision too. The first model was Image Transformer (Parmar et al., 2018). It restricts the self-attention mechanism to attend to local neighborhoods, so larger images can be processed. Class-conditioned generation is also supported, by passing the embedding of the relative class in input. To avoid restricting self-attention to local neighborhoods, Vision Transformer (Dosovitskiy et al., 2021) divides an image into fixed-size patches, linearly embeds each of them, adds position embeddings, and then feeds the resulting sequence of vectors to a standard Transformer encoder. Masked Autoencoders (MAE) (He et al., 2022) instead uses an encoder-decoder architecture based on Transformers trained with masked image modeling (i.e., to reconstruct randomly masked pixels). A BERT adaptation to images called Bidirectional Encoder representation from Image Transformers (BEiT) (Bao et al., 2022) has also been proposed. Masked image modeling has also been used together with classic autoregressive loss (Chen et al., 2020). Conversely, Vector Quantised-GAN (VQ-GAN) (Esser et al., 2021) allows a Transformer to be based on vector quantization. A GAN learns an effective codebook of image constituents. To do so, the generator is implemented as an auto-encoder; vector quantization is applied over the latent representation returned by the encoder. It is then possible to efficiently encode an image in a sequence corresponding to the codebook-indices of their embeddings. The Transformer is finally trained on that sequence to learn long-range interactions. These changes also allow us to avoid quadratic scaling, which is intractable for high-resolution images. Finally, DALL-E (Radford et al., 2021) takes advantage from a discrete VAE. To generate images based on an input text, it learns a discrete image encoding; it concatenates the input text embedding with the image encoding; it learns autoregressively on them. CogView implements a similar architecture (Ding et al., 2021).
@@
Finally, Transformer-based models have also been used in multimodal settings, in which data sources are of different types. A survey can be found in (Suzuki and Matsuo, 2022). The first examples of these systems consider text and images together as output of the Transformer architecture. By aligning their latent representations, images and texts can be generated by Transformer-based decoders given a multimodal representation. For instance, Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021) has an image encoder pre-trained together with a text encoder to generate a caption for an image. A Large-scale ImaGe and Noisy-text embedding (ALIGN) (Jia et al., 2021), based on similar mechanisms, can achieve remarkable performance through training based on a noisier dataset. In (Tsimpoukelli et al., 2021) the authors propose a frozen language model for multimodal few-shot learning: a vision encoder is trained to represent each image as a sequence of continuous embeddings, so that the frozen language model prompted with this embedding can generate the appropriate caption. In (Fei et al., 2022) the authors present Bridging-Vision-and-Language (BriVL), which performs multimodal tasks by learning from weak semantic correlation data. Finally, there is a trend toward even more complex multimodal models. For example, Video-Audio-Text Transformer (VATT) (Akbari et al., 2021) learns to extract multimodal representations from video, audio and text; instead, Gato (Reed et al., 2022) serializes all data (e.g., text, images, games, other RL-related tasks) into a flat sequence of tokens that is then embedded and passed to a standard large-scale language model. Similarly, Gemini (Gemini Team and Google, 2023) achieves state-of-the-art performance in multimodal tasks by working on interleaved sequences of text, image, audio, and video as inputs; (Gemini Team and Google, 2024) extends it to Mixture-of-Experts setting. Finally, NExT-GPT (Wu et al., 2023) handles any combination of four modalities (text, audio, image, and video) by connecting a language model with multimodal adaptors and diffusion decoders (see Section 3.5).
@@ 3.4.3. Applications
Transformer-based large language models can be used for almost any NLP task, including text summarization, generation and interaction. In order to do so, the model can be used as frozen (i.e., to provide latent representations in input to other models); can be fine-tuned for the specific objective; can be exploited with zero-shot, one-shot or few-shot setting by prompting the task or few demonstrations in input. Transfer learning can instead be used to perform image classification by means of Transformer-based models trained on images. Other domain-specific techniques can be used as well: for instance, PlotMachines (Rashkin et al., 2020) learns to write narrative paragraphs not by receiving prompts, but by receiving plot outlines and representations of previous paragraphs.
From a generative learning perspective, Transformers have shown impressive performance in producing long sequences of texts and music or speech (Wang et al., 2023), as well as in generating images based on input text. It is worth noting that their application has not been limited to these data sources. For instance, AlphaFold uses a Transformer architecture to predict protein structure (Jumper et al., 2021); RecipeGPT employs it to generate recipes (Lee et al., 2020); and GitHub Copilot relies on it to support code development (Chen et al., 2021).
@@ 3.4.4. Critical Discussion
Given the fact that the Transformers can be considered as an evolution of sequence prediction models, the observations made for that class of models (see Section 3.3) apply also to them. However, the inherent characteristics of their architecture allow for larger models and higher-quality outputs, which are also able to capture a variety of dependencies of text and across data sources. More in general, a broader conceptual space is induced. This means that domain-specific tasks might be addressed by means of solutions outside or at the boundary of the sub-space linked with that domain. Moreover, possibly also through a careful use of inputs (see Section 3.7), their adoption might lead to transformational creativity. As far as Boden’s criteria are concerned, there is no guarantee that the output of the Transformer architecture would be valuable, novel, or surprising, even though current state-of-the-art large language models (LLMs) achieve almost human-like performance in creative tests (Stevenson et al., 2022; Zhao et al., 2024).
@@ 3.5.1. Core Concepts
Diffusion models are a family of methods able to generate samples by gradually removing noise from a signal (Sohl-Dickstein et al., 2015). The most representative approach is the Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020). An input $\mathbf{x_{0}}$ is corrupted by gradually adding noise until obtaining an $\mathbf{x_{T}}$ from a pre-defined distribution; the model then has to reverse the process. Each timestep $t$ corresponds to a certain noise level; $\mathbf{x_{t}}$ can be seen as a mixture of $\mathbf{x_{0}}$ with some noise $\mathbf{\epsilon}$ whose ratio is determined by $t$. The model learns a function $\epsilon_{\theta}$ to predict the noise component of $\mathbf{x_{t}}$ by minimizing the mean-squared error. $\mathbf{x_{t-1}}$ is then obtained from a diagonal Gaussian with mean as a function of $\epsilon_{\theta}\!\left(\mathbf{x_{t}},t\right)$, and with a fixed (Ho et al., 2020) or learned (Nichol and Dhariwal, 2021) variance. In other words, it learns to associate points from a predefined random distribution with real data through iterative denoising. Because of it, at inference time, a diffusion model can generate a new sample by starting from pure random noise. The generation can also be conditioned by simply modifying the noise perturbation so that it depends on the conditional information. However, this iterative sampling process might potentially lead to slow generation; a proposed solution is to induce self-consistency, i.e., ensuring that points on the same trajectory map to the same initial ones (Song et al., 2023). In this way, the output can be obtained in a single step.

### v6

@@ 3.2.3. Applications
GANs have been applied to a variety of practical problems in several application scenarios. They have been widely used for semi-supervised learning (Odena, 2016); for generating adversarial examples (Xiao et al., 2018) to better train image classifiers (Madry et al., 2018); and, in general, in computer vision (see (Wang et al., 2021) for a detailed discussion). The generative power of GANs has also found its place in recommender systems (see (Deldjoo et al., 2021)) to generate fashion items; in science and chemistry (Mendez-Lucio et al., 2020; Motamed et al., 2021). Of course, its ability to generate high-quality samples has been exploited in many other areas, from anime design (Jin et al., 2017) and 3D object modeling (Wu et al., 2016) to photo-realistic consequences of climate change (Schmidt et al., 2022). Conditional inputs also allow the production of artistic works by controlling stylistic properties such as genre (Tan et al., 2017) or influencer (Chu, 2018). Finally, the most famous example of the artistic power of GAN is the collection of paintings by Obvious, a French art collective (Vernier et al., 2020); one of their works has been sold to more than 400,000 dollars [^3] Fun fact: the sold painting is called Portrait of Edmond De Belamy because Belamy sounds like bel ami, a sort of French translation of… Goodfellow..
@@ 3.3.1. Core Concepts
A sequence prediction model is a generative model that considers generation as a sequential process. It works in an autoregressive fashion: it predicts the future outcome of the sequence (i.e., the next token) from the previously observed outcomes of that sequence, usually by means of an internal state that encodes information from the past. It is trained to minimize the prediction error for each token in the dataset. At inference time, this simple yet effective approach only requires to produce one token after the other, feeding back to the model what has been produced so far (Karpathy, 2015). It makes it possible to learn dependencies between tokens in real data so that the same dependencies can be exploited when generating synthetic data. However, this causes the generation to be highly dependent on real data, e.g., there is the risk of potentially reproducing portions of the training set.
@@ 3.3.2. Examples of Models
Sequence prediction models are also used for domains not commonly modeled as sequences, like images. Image modeling can be defined in a discrete way by means of a joint distribution of pixels: the model learns to predict the next pixel given all the previously generated ones. It starts at the top left pixel and then proceeds towards the bottom right. The two seminal architectures for sequence prediction of images are PixelRNN and PixelCNN (Van Den Oord et al., 2016b). The former is a two-dimensional RNN (based on rows or diagonals). The latter is a convolutional neural network (CNN) with an additional fixed dependency range (i.e., the convolution filters are masked in order to only use information about pixels above and to the left of the current one). To obtain better results, gated activation units can be used in place of rectified linear units between the masked convolutions; conditional inputs encoding high-level image descriptions can be used as well (Van Den Oord et al., 2016c). Notably, the Gated PixelCNN architecture can also be used for other types of data: WaveNet (Van Den Oord et al., 2016a) implements it to generate audio based on the waveform, possibly guiding the generation with conditional inputs.
@@ 3.3.3. Applications
As discussed, sequence prediction models have been used to learn to write poems or stories (by predicting a character, syllable, or word after the other); to compose music (by predicting a note or a waveform after the other); to draw images (by predicting a pixel after the other). In general, they can be used for any kind of time series forecasting (Lim and Zohren, 2021). They can also be used for co-creativity, as in Creative Help (Roemmele and Gordon, 2018). Despite their simplicity, sequence prediction models are one of the most successful generative techniques. An interesting example is Sunspring. It might be considered as the first AI-scripted movie: it was generated by a Char-RNN trained on thousands of sci-fi scripts (Miller, 2019). The quality of the result is demonstrated by the fact that it was able to reach the top ten at the annual Sci-Fi London Film Festival in its 48-Hour Film Challenge [^4] Quite interestingly, the AI that wrote Sunspring declared that its name was Benjamin, probably in honor of Walter Benjamin, the German philosopher who, already in 1935 (Benjamin, 2008), understood that new mechanical techniques related to art can radically change the public attitude to art and artists..
@@ 3.3.4. Critical Discussion
Sequence prediction models generate outputs that have characteristics of both exploratory and combinatorial creativity. They are based on probabilistic predictions and they are able to generate new outputs in the induced space, but they can also reuse sequences of tokens from different works, combining them together. There is no guarantee that the results will be valuable or novel, and classic methods such as RNNs lack surprise (Bunescu and Uduehi, 2019). It is worth noting that the possibility of using conditional inputs and being able to work at different levels of abstraction might indirectly lead to creative outputs, but creativity should then be attributed to the higher-level component (or human if the input is provided by the user) that is guiding the generation for specific elements and characteristics of the result.
@@ 3.4.1. Core Concepts
In principle, a Transformer is nothing more than an autoregressive model: it works by predicting the current token given the previous ones (see Section 3.3). However, few fundamental differences exist. A Transformer can also be trained by means of masked modeling: some of the input tokens are randomly masked, and the model has to learn how to reconstruct them from the entire context, and not only from the previous portions (Devlin et al., 2019). The possibility of dealing with very long sequences allows for prompting. By providing a natural language prompt in input, the model can generate the desired output, e.g., the answer to a question, a class between a set of classes for a given text, or a poem in a particular style (Brown et al., 2020). This is done by simply passing the prompt in input as a text, and then leveraging the model to predict what comes next (e.g., the answer to a question). These advantages, together with the very large amount of data available, the increasing computational power, and the parallelism induced by their architecture, have led Transformer-based models to become the state of the art for several tasks. Nevertheless, the computational costs of the architecture from (Vaswani et al., 2017) grow quadratically with the input size.
@@ 3.4.2. Examples of Models
The domain mostly influenced by Transformers is natural language processing (NLP). Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2019) is a Transformer-based encoder trained for both predicting the next sentence (in an autoregressive fashion) and reconstructing masked tokens from the context. Several enhanced variations of the original model have been proposed, such as, for instance, solutions that remove the next-sentence pre-training objective (Liu et al., 2019), use inter-sentence coherence as an additional loss (Lan et al., 2020), or employ distillation (Hinton et al., 2015) to train a smaller model (Sanh et al., 2019). The other main approach is that used by the Generative Pre-trained Transformer (GPT) family (Radford, 2018; Radford et al., 2019; Brown et al., 2020). Here, a Transformer-based decoder is trained in an autoregressive way by additionally conditioning on the task of interest. After training, it can be used to perform a wide range of tasks by providing a description or a few demonstrations of the task. The effectiveness of this text-to-text generative approach has then been explored by T5 (Raffel et al., 2020). Many other large language models (Shoeybi et al., 2019; Touvron et al., 2023a; Zhang et al., 2022) have been proposed to achieve better results by means of more parameters and computation (Smith et al., 2022), or more qualitative data (Gunasekar et al., 2023). Mixture of Experts (Shazeer et al., 2017) can be used as well in place of the feed-forward network to train a larger but lighter model (since only portions of it are used per task), as done by Generalist Language Model (GLaM) (Du et al., 2022). Finally, Bidirectional and Auto-Regressive Transformer (BART) (Lewis et al., 2020) ideally merges together a BERT-encoder (trained by corrupting text with an arbitrary noising function) and a GPT-decoder (trained to reconstruct the original text autoregressively). Such an encoder-decoder architecture is able to achieve state-of-the-art results in machine translations, as well as other text-to-text tasks.
@@
Transformers have been highly influential in computer vision too. The first model was Image Transformer (Parmar et al., 2018). It restricts the self-attention mechanism to attend to local neighborhoods, so larger images can be processed. Class-conditioned generation is also supported, by passing the embedding of the relative class in input. To avoid restricting self-attention to local neighborhoods, Vision Transformer (Dosovitskiy et al., 2021) divides an image into fixed-size patches, linearly embeds each of them, adds position embeddings, and then feeds the resulting sequence of vectors to a standard Transformer encoder. Masked Autoencoders (MAE) (He et al., 2022) instead uses an encoder-decoder architecture based on Transformers trained with masked image modeling (i.e., to reconstruct randomly masked pixels). A BERT adaptation to images called Bidirectional Encoder representation from Image Transformers (BEiT) (Bao et al., 2022) has also been proposed. Masked image modeling has also been used together with classic autoregressive loss (Chen et al., 2020). Conversely, Vector Quantised-GAN (VQ-GAN) (Esser et al., 2021) allows a Transformer to be based on vector quantization. A GAN learns an effective codebook of image constituents. To do so, the generator is implemented as an auto-encoder; vector quantization is applied over the latent representation returned by the encoder. It is then possible to efficiently encode an image in a sequence corresponding to the codebook indices of their embeddings. The Transformer is finally trained on that sequence to learn long-range interactions. These changes also allow us to avoid quadratic scaling, which is intractable for high-resolution images. Finally, DALL-E (Radford et al., 2021) takes advantage of a discrete VAE. To generate images based on an input text, it learns a discrete image encoding; it concatenates the input text embedding with the image encoding; it learns autoregressively on them. CogView implements a similar architecture (Ding et al., 2021).
@@
Finally, Transformer-based models have also been used in multimodal settings, in which data sources are of different types. A survey can be found in (Suzuki and Matsuo, 2022). The first examples of these systems consider text and images together as the output of the Transformer architecture. By aligning their latent representations, images and texts can be generated by Transformer-based decoders given a multimodal representation. For instance, Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021) has an image encoder pre-trained together with a text encoder to generate a caption for an image. A Large-scale ImaGe and Noisy-text embedding (ALIGN) (Jia et al., 2021), based on similar mechanisms, can achieve remarkable performance through training based on a noisier dataset. In (Tsimpoukelli et al., 2021) the authors propose a frozen language model for multimodal few-shot learning: a vision encoder is trained to represent each image as a sequence of continuous embeddings, so that the frozen language model prompted with this embedding can generate the appropriate caption. In (Fei et al., 2022) the authors present Bridging-Vision-and-Language (BriVL), which performs multimodal tasks by learning from weak semantic correlation data. Finally, there is a trend toward even more complex multimodal models. For example, Video-Audio-Text Transformer (VATT) (Akbari et al., 2021) learns to extract multimodal representations from video, audio, and text; instead, Gato (Reed et al., 2022) serializes all data (e.g., text, images, games, other RL-related tasks) into a flat sequence of tokens that is then embedded and passed to a standard large-scale language model. Similarly, Gemini (Gemini Team and Google, 2023) achieves state-of-the-art performance in multimodal tasks by working on interleaved sequences of text, image, audio, and video as inputs; (Gemini Team and Google, 2024) extends it to Mixture-of-Experts setting. Finally, NExT-GPT (Wu et al., 2023) handles any combination of four modalities (text, audio, image, and video) by connecting a language model with multimodal adaptors and diffusion decoders (see Section 3.5).
@@ 3.4.3. Applications
Transformer-based large language models can be used for almost any NLP task, including text summarization, generation, and interaction. In order to do so, the model can be used as frozen (i.e., to provide latent representations in input to other models); can be fine-tuned for the specific objective; can be exploited with zero-shot, one-shot or few-shot setting by prompting the task or few demonstrations in input. Transfer learning can instead be used to perform image classification by means of Transformer-based models trained on images. Other domain-specific techniques can be used as well: for instance, PlotMachines (Rashkin et al., 2020) learns to write narrative paragraphs not by receiving prompts, but by receiving plot outlines and representations of previous paragraphs. From a generative learning perspective, Transformers have shown impressive performance in producing long sequences of texts and music or speech (Wang et al., 2023), as well as in generating images based on input text. Their application has not been limited to these data sources. For instance, AlphaFold uses a Transformer architecture to predict protein structure (Jumper et al., 2021); RecipeGPT employs it to generate recipes (Lee et al., 2020); and GitHub Copilot relies on it to support code development (Chen et al., 2021).
@@ 3.4.4. Critical Discussion
Given the fact that the Transformers can be considered as an evolution of sequence prediction models, the observations made for that class of models (see Section 3.3) apply also to them. However, the inherent characteristics of their architecture allow for larger models and higher-quality outputs, which are also able to capture a variety of dependencies of text across data sources. More in general, a broader conceptual space is induced. This means that domain-specific tasks might be addressed by means of solutions outside or at the boundary of the sub-space linked with that domain. Moreover, possibly also through careful use of inputs (see Section 3.7), their adoption might lead to transformational creativity. As far as Boden’s criteria are concerned, there is no guarantee that the output of the Transformer architecture would be valuable, novel, or surprising, even though current state-of-the-art large language models (LLMs) achieve almost human-like performance in creative tests (Stevenson et al., 2022; Zhao et al., 2024).
@@ 3.5.1. Core Concepts
Diffusion models are a family of methods able to generate samples by gradually removing noise from a signal (Sohl-Dickstein et al., 2015). The most representative approach is the Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020). An input $\mathbf{x_{0}}$ is corrupted by gradually adding noise until obtaining an $\mathbf{x_{T}}$ from a pre-defined distribution; the model then has to reverse the process. Each timestep $t$ corresponds to a certain noise level; $\mathbf{x_{t}}$ can be seen as a mixture of $\mathbf{x_{0}}$ with some noise $\mathbf{\epsilon}$ whose ratio is determined by $t$. The model learns a function $\epsilon_{\theta}$ to predict the noise component of $\mathbf{x_{t}}$ by minimizing the mean-squared error. $\mathbf{x_{t-1}}$ is then obtained from a diagonal Gaussian with mean as a function of $\epsilon_{\theta}\!\left(\mathbf{x_{t}},t\right)$, and with a fixed (Ho et al., 2020) or learned (Nichol and Dhariwal, 2021) variance. In other words, it learns to associate points from a predefined random distribution with real data through iterative denoising. Because of this, at inference time, a diffusion model can generate a new sample by starting from pure random noise. The generation can also be conditioned by simply modifying the noise perturbation so that it depends on the conditional information. However, this iterative sampling process might potentially lead to slow generation; a proposed solution is to induce self-consistency, i.e., ensuring that points on the same trajectory map to the same initial ones (Song et al., 2023). In this way, the output can be obtained in a single step.